# CodeEditorBench: Evaluating Code Editing Capability of Large Language Models

### v2

@@ Abstract
${}^{\ast}$ These authors contributed ${}^{{\dagger}}$ Corresponding Authors.
@@ 3.1 Problem Definition
  • an ideal debugger D that rectifies any buggy code from $c$ to $*c*$ should satisfy that D($c$) = $*c*$ s.t.$\forall$($x_{i}$, $y_{i}$) $\in R_{i}$, $*a_{c}*$($x_{i}$) = $y_{i}$. Debugging can be regarded as the converting process of debugger D.
  • an ideal translater T that translates any code from language a to language b should satisfy that T($c$) = $*c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{i}$, $a_{c}$($x_{i}$) = $y_{i}$ and $*a_{c}*$($x_{i}$) = $y_{i}$. Translating can be regarded as the converting process of translater T.
  • an ideal optimizer P that make any code more effective in either time complexity or space complexity with the guarantee of accuracy should satisfy that P($c$) = $*c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{i}$, $a_{c}$($x_{i}$) = $y_{i}$,$*a_{c}*$($x_{i}$) = $y_{i}$ and $avg_{time}(c*)\leq avg_{time}(c)$ or $avg_{memory}(c*)\leq avg_{memory}(c)$. Polishing can be regarded as the converting process of optimizer P
  • an ideal requirement switcher S that switches a similar code that satisfies question $a$ to a target code that satisfies question $b$ based on the sample inputs and outputs of $a$ and $b$ should satisfy that S($c$) = $*c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{b}$, $*a_{c}*$($x_{i}$) = $y_{i}$.Switching can be regarded as the converting process of switcher S.
@@ Appendix B Evaluated Models
ise-uiuc/Magicoder-S-DS-6.7B
6.7B 2023-12-04
Magicoder-S-DS-6.7B
ise-uiuc/Magicoder-S-CL-7B
7B 2023-12-04
Magicoder-S-CL-7B
bigcode/octocoder
15.5B 2023-08-14
octocoder
WizardLM/WizardCoder-15B-V1.0
15B 2023-06-16
WizardCoder-15B-V1.0
WizardLM/WizardCoder-33B-V1.1
33B 2024-01-04
WizardCoder-33B-V1.1
deepseek-ai/deepseek-coder-33b-instruct
33B 2023-02-01
deepseek-coder-33b-instruct
codefuse-ai/CodeFuse-CodeLlama-34B
34B 2023-09-11
CodeFuse-CodeLlama-34B
Phind/Phind-CodeLlama-34B-v2
34B 2023-09-01
Phind-CodeLlama-34B-v2
m-a-p/OpenCodeInterpreter-DS-6.7B
6.7B 2024-02-22
OpenCodeInterpreter-DS-6.7B
m-a-p/OpenCodeInterpreter-DS-33B
33B 2024-02-22
OpenCodeInterpreter-DS-33B
codellama/CodeLlama-34b-hf
34B 2023-08-24
CodeLlama-34b-hf
codellama/CodeLlama-7b-Instruct-hf
7B 2023-08-24
CodeLlama-7b-Instruct-hf
codellama/CodeLlama-13b-Instruct-hf
13B 2023-08-24
CodeLlama-13b-Instruct-hf
codellama/CodeLlama-34b-Instruct-hf
34B 2023-08-24
CodeLlama-34b-Instruct-hf
gpt-3.5-turbo-1106 (OpenAI, 2022)
- 2021-10-01
gpt-3.5-turbo-1106
gpt-4-0613 (OpenAI, 2023)
- 2021-10-01
gpt-4-0613
glm-4 (zhipuai, 2024)
- 2024-01-16
glm-4
gemini-pro (Google, 2023)
- 2023-12-06
gemini-pro
gemini-ultra (Google, 2023)
- 2023-12-06
@@ Appendix E Pass Criteria
Polish $\bar{T}$ms $\bar{M}$
Pass all test cases
and
$\bar{T}_{\text{avg}}<\bar{T}$ or $\bar{M}_{\text{avg}}<\bar{M}$

### v3

@@ Abstract
These authors contributed equally. Corresponding Authors.
@@ 3.1 Problem Definition
  • an ideal debugger D that rectifies any buggy code from $c$ to $c*$ should satisfy that D($c$) = $c*$ s.t.$\forall$($x_{i}$, $y_{i}$) $\in R_{i}$, $a_{c}*$($x_{i}$) = $y_{i}$. Debugging can be regarded as the converting process of debugger D.
  • an ideal translater T that translates any code from language a to language b should satisfy that T($c$) = $c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{i}$, $a_{c}$($x_{i}$) = $y_{i}$ and $a_{c}*$($x_{i}$) = $y_{i}$. Translating can be regarded as the converting process of translater T.
  • an ideal optimizer P that make any code more effective in either time complexity or space complexity with the guarantee of accuracy should satisfy that P($c$) = $c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{i}$, $a_{c}$($x_{i}$) = $y_{i}$,$a_{c}*$($x_{i}$) = $y_{i}$ and $avg_{time}(c*)\leq avg_{time}(c)$ or $avg_{memory}(c*)\leq avg_{memory}(c)$. Polishing can be regarded as the converting process of optimizer P
  • an ideal requirement switcher S that switches a similar code that satisfies question $a$ to a target code that satisfies question $b$ based on the sample inputs and outputs of $a$ and $b$ should satisfy that S($c$) = $c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{b}$, $a_{c}*$($x_{i}$) = $y_{i}$.Switching can be regarded as the converting process of switcher S.
@@ Appendix B Evaluated Models
ise-uiuc/Magicoder-S-DS-6.7B 6.7B 2023-12-04 Magicoder-S-DS-6.7B
ise-uiuc/Magicoder-S-CL-7B 7B 2023-12-04 Magicoder-S-CL-7B
bigcode/octocoder 15.5B 2023-08-14 octocoder
WizardLM/WizardCoder-15B-V1.0 15B 2023-06-16 WizardCoder-15B-V1.0
WizardLM/WizardCoder-33B-V1.1 33B 2024-01-04 WizardCoder-33B-V1.1
deepseek-ai/deepseek-coder-33b-instruct 33B 2023-02-01 deepseek-coder-33b-instruct
codefuse-ai/CodeFuse-CodeLlama-34B 34B 2023-09-11 CodeFuse-CodeLlama-34B
Phind/Phind-CodeLlama-34B-v2 34B 2023-09-01 Phind-CodeLlama-34B-v2
m-a-p/OpenCodeInterpreter-DS-6.7B 6.7B 2024-02-22 OpenCodeInterpreter-DS-6.7B
m-a-p/OpenCodeInterpreter-DS-33B 33B 2024-02-22 OpenCodeInterpreter-DS-33B
codellama/CodeLlama-34b-hf 34B 2023-08-24 CodeLlama-34b-hf
codellama/CodeLlama-7b-Instruct-hf 7B 2023-08-24 CodeLlama-7b-Instruct-hf
codellama/CodeLlama-13b-Instruct-hf 13B 2023-08-24 CodeLlama-13b-Instruct-hf
codellama/CodeLlama-34b-Instruct-hf 34B 2023-08-24 CodeLlama-34b-Instruct-hf
gpt-3.5-turbo-1106 (OpenAI, 2022) - 2021-10-01 gpt-3.5-turbo-1106
gpt-4-0613 (OpenAI, 2023) - 2021-10-01 gpt-4-0613
glm-4 (zhipuai, 2024) - 2024-01-16 glm-4
gemini-pro (Google, 2023) - 2023-12-06 gemini-pro
gemini-ultra (Google, 2023) - 2023-12-06 -
@@ Appendix E Pass Criteria
Polish $\bar{T}$ms $\bar{M}$ Pass all test cases and $\bar{T}_{\text{avg}}<\bar{T}$ or $\bar{M}_{\text{avg}}<\bar{M}$